OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. - tan^2x + sec^2x = 1

OpenStudy (raden):

first, use these identity : - tan^2x = - sin^2 x/cos^2 x sec^2 x = 1/cos^2 x see ur equation can becomes - tan^2x + sec^2x = - sin^2 x/cos^2 x + 1/cos^2 x = (- sin^2 x + 1)/cos^2 c remember the identity that sin^2 x + cos^2 x = 1 or -sin^2 x + 1 = cos^2 x so, it can be = cos^2 x/cos^2 x = 1 LHS = RHS

OpenStudy (raden):

another way is if u remember the identity ; sec^2 x = 1 + tan^2 x so, - tan^2x + sec^2x = - tan^2x + 1 + tan^2 x = 1 that's easier than before